### v3

@@ 5.3.2 Trace Modules
ZEKRA [36] implicitly relies on a trusted trace module with a certified asymmetric key pair in order to provide path authenticity. The proposal does not specify an implementation per se, but specifies that the module must chronologically and faithfully provide a trace of the executed path when executing a program region to be attested.
@@ 5.3.3 Performance Monitoring Units
LightFAt relies on a CPU’s PMU on $\mathcal{P}$ to monitor the number of instructions per cycles and L1 cache accesses used during the execution of the target program for a given challenge from $\mathcal{V}$. The number of events is sent to $\mathcal{V}$ who then uses a machine learning model to classify the program execution as being malicious or legitimate. The model is assumed to be trained on the target program pre-deployment with known, labelled malicious/legitimate traces. Results of 95% accuracy and 0.9757 F1-score are presented using a AMD Ryzen 7 2700X 64-bit processor.
5.4 Discrete Hardware Modules
@@
Geden & Rasmussen [47] proposed a hardware security module (HSM) that collects control-flow information from the system bus (Figure 14). The HSM continuously takes control-flow events as input from the CPU, including the requested instruction address, opcode, and operand values over the system-on-chip’s control line and address and data buses. The collected events are compared with security policies defined in the HSM’s runtime integrity model for detecting code injection and re-use attacks. The use of continuous run-time data addresses TOCTOU attacks: a challenge-response protocol that measures events at only fixed points in time does not detect attackers that act transiently between those points. The proposal allows $\mathcal{V}$ to acquire a report from the HSM showing whether any security policy violations were detected; it also supports local verification whereby the module outputs a binary result to indicate if a target control-flow hijacking attack is detected. In scenarios where attacks are detected, the module outputs a binary result and provides auxiliary information, such as the last executed instruction.
@@
Tiny-CFA [62] uses a minimalist approach to control-flow attestation, focussing on reducing the trusted computing base and the number of invasive modifications needed to realise a trust anchor. The work employs a simple external hardware module that monitors run-time control-flow states in order to produce a proof-of-execution (PoX). Tiny-CFA proposes an approach that combines automated code instrumentation with APEX [74], which uses a bespoke hardware module to provide a PoX from a micro-controller unit core.
@@ 6.1 Compiler-assisted Instrumentation
Where $cp_{A}$ and $cp_{B}$ are checkpoint identifiers representing an arbitrary sequence of BBLs. $H(LoA)$ is the hashed measurement value of a list of actions (LoA), which is the list of CFG edges for traversing $cp_{A}\to cp_{B}$; and $(BBL_{si},BBL_{di})$ is the source and destination BBL at index $i$. The compiler is also used for detecting and instrument control-flow instances, which are passed to three kernel-space modules responsible for generating measurements. To support this, the Measurement Generator analyses the LLVM intermediate representation (IR) of the target program’s processed source code. The CRAB framework is used to generate the CFG itself, which is used to resolve indirect forward jumps. We note that mapping LLVM to raw assembly BBLs involves the use of dummy code, which is later removed using a custom binary rewriting tool. The authors add approximately 3.5K lines of code (LOC) to CRAB and LLVM 5.0.
@@
LAPE [49] uses the LLVM-compatible ACES [79] compiler to create and instrument ‘compartments’ on ARMv7 devices. Programs execute within components where indirect calls and function entry and exit points are recorded and measured using a hash-based approach. We note that, despite LLVM’s flexibility, there are some downsides. LLVM is typically not used unchanged: proposals tend to require extensive modifications or extensions of the magnitude of 1,000s LOC [10, 13, 14]. In practice, this may pose compatibility issues when adapting the schemes to existing toolchains and development environments.
In other work, Hu et al. [11] use a combination of compiler-assisted instrumentation and binary rewriting—discussed in §6.2—to hook control-flow instances. The GCC compiler with the -finstrument-functions flag is used to instrument function entry and exit points. Afterwards, branches within functions are hooked using binary rewriting using the Capstone disassembly framework [80].
@@ 6.2 Binary Analysis and Rewriting
Binary rewriting involves making direct modifications to the compiled target program, where control-flow instances are represented and changed in the form of machine code. This approach can be applied to existing programs, making it suitable for incorporating CFA into software where the original source code is not accessible. The approach has particular applications to retrospectively protecting legacy and proprietary software.
@@
Recording every function call can impose significant performance overhead (see p. 3, [58]). To address this, ReCFA [58] derives a CFG from the target program using TypeArmor for performing call-site filtering, whereby CFG nodes are skipped where none of its predecessors has more than one successor. ReCFA uses Dyninst [81] for the static binary instrumentation of user-space programs on $\mathcal{P}$. Dyninst provides a set of libraries and utilities designed for the static (and dynamic) inspection and manipulation of executables. Dyninst enables users to insert, modify, or remove code in a program’s CFG without access to its source code. Intel’s MPK system is used to protect the integrity of instrumented data structures—the loop stack and path stack—used by control-flow events. The authors compare this to SCaRR [13] where control-flow edges are measured in kernel space, which has a “potential enormous context switching cost for the control-flow events on an order of magnitude in gigabytes” (p. 6, [58]). The authors combine kernel mode-based functions to preserve code integrity and MPK-protect data structures in order to uphold the integrity of the measurement procedure.
@@ 7.2 Other Path Log Representations
Another example is RAGE [42], which relies on a machine learning-based approach using unsupervised graph neural networks to avoid computing a complete CFG of the target program. In the training phase, $\mathcal{V}$ uses one legitimate control-flow path represented as a sequence of BBL addresses as a partial CFG (PCFG) with which feature extraction is performed. The features are used to train a Variational Graph Autoencoder (VGAE) model [88] that encodes CFGs into a lower-dimensional space (graph embedding). The legitimate control-flow path is converted to an embedding which serves as a reference. Ten other legitimate and illicit control-flow paths are used to set an appropriate threshold using the directed Hausdorff distance [89] to quantify the difference between two embeddings. During the attestation process, $\mathcal{P}$ sends a measurement result, a sequence of BBL addresses, executed by the target program to $\mathcal{V}$. $\mathcal{V}$ inputs this data to the VGAE model to produce a new embedding. $\mathcal{V}$ compares the embedding with the referenced embedding; if their distance exceeds a predefined threshold, then it indicates the detection of attacks. The authors tested their proposal using synthetic ROP and DOP attacks, yielding results of 0.9803 (ROP) and 0.9101 (DOP) F1-score while maintaining a low False Positive Rate of 0.0319.
@@ 7.5 Hybrid Measurement
We define hybrid measurement as combining more than one of the measurement methods above. Currently, three existing CFA proposals fall into the category, which combine hash-based measurement and abstract execution. LO-FAT [12] uses abstract execution for loop paths and hash-based measurement for the rest of the control-flow path; avoiding hash-based measurements for loops avoids the path explosion problem. OAT [6] is the first CFA scheme to explicitly address control-flow and data-only attacks on embedded systems. OAT addresses this through a combination of hash-based measurements and compact execution traces using software-only techniques in a TEE (ARM TrustZone).
@@ 8 Evaluation
G&R [47] 2019 Hardware Module $\blacktriangle$ $\blacktriangle$ Conventional (C.2)(C.3)(C.4)(C.7) (A.1)(MP)(A.2)(A.3)(A.4)
Liu et al. [7] 2019 TEE (ARM TrustZone) + PUF Whitelist Binary Rewriting Conventional (C.2)(C.3)(C.5)(C.6)(C.7) (A.1)(A.2)(A.3)
@@
[…]
@@
Ben Yehuda et al. [39] 2022 Kernel Hash-based Binary Rewriting Conventional (C.1) (A.1)(A.2)(A.3)
@@
GuaranTEE [10] 2022 TEE (Intel SGX) Abstract Execution Compiler (LLVM) Hybrid (C.4)(C.5)(C.6)(C.7) (A.1)(A.2)
@@
ARI [43] 2023 TEE (ARM TrustZone) Abstract Execution Compiler (LLVM) Conventional (C.8) (A.1)(A.2)(A.3)
@@
ZEKRA [36] 2023 Trace Module* Hash-based $\blacktriangle$ * VC (C.1) (A.1)(A.2)
RAGE [42] 2024 TEE* Path Log Dynamic Conventional (C.1) (A.2)$\pm$(A.4)$\pm$
LightFAt [5] 2024 PMU Path Log Dynamic Conventional (C.1) (A.1)$\pm$(A.2)$\pm$
@@ 8.1 Comparison Criteria
  • $\mathcal{P}$-$\mathcal{V}$ paradigm: We give the prover-verifier ($\mathcal{P}$-$\mathcal{V}$) paradigm used by the proposal; for example, whether it uses a conventional remote attestation, hybrid, or another approach described in §4.
@@ 8.2 Discussion and Analysis
Many CFA schemes claimed the ability to detect code injection attacks, i.e. (A.1). However, a significant minority of proposals (7/30, $\approx$23%) do not provide this due to features bestowed by the scheme itself. Rather, they make assumptions regarding the availability of mechanisms that enforce some aspect of the target program’s immutability; for example, using executable-space, data execution prevention, and other memory protection mechanisms [12, 47]. LO-FAT, for instance, assumes that “the adversary cannot modify program code at run-time (marked as rx)” (p. 3, [12]) as an implicit requirement for a memory protection mechanism. Geden & Rasmussen [47] requires non-writable code regions and non-executable data regions. In a TEE-based scheme, OAT [6] assumes that the attackers cannot tamper with the instrumented code or the trampoline library in the normal world, citing some memory protection methods for embedded devices as potential solutions [92, 93]. [^7] We note that, in some cases, it is possible to bypass memory protection mechanisms, e.g. DEP, using ROP-style attacks. To this end, we refer the reader to work by Göktas et al. [94].
We observe that ROP attacks are the primary threat in schemes that address code reuse attacks, (A.2). Such proposals add detection mechanisms to control-flow transfers that return from functions. Some work, e.g. C-FLAT [4] and GuaranTEE [10], address a particular variant in which a function is invoked from multiple calling locations. Suppose there are two legitimate control-flow paths, $A\to B\to C$ and $D\to B\to E$, where $B$ is a vulnerable function. An attacker may perform an attack to alter the control-flow path to $D\to B\to C$, which still visits legitimate CFG nodes and may not be detected as an illicit path. This has been addressed through the use of shadow stacks [47, 58, 7, 13]. When a call-based control-flow transfer is executed, if its target address is valid, then its call-after point will be pushed on this shadow stack. For a return-based transfer, if its target address is valid based on CFG, then $\mathcal{V}$ checks if it is identical to the top element of the shadow stack. If it is not, a control-flow hijacking attack would be detected. If it matches, the top element is popped from the stack. Call-return matching is another solution in which $\mathcal{V}$ marks legitimate pairs of call and return control-flow transfers within the scheme’s offline phase, which are measured at run-time [4].
@@
A final point of note is the relative lack of focus on direct data manipulation (DDM) and data-oriented programming (DOP) attacks: (A.3) and (A.4) respectively. 16/31 ($\approx$52%) and only 5/31 ($\approx$16%) proposals address (A.3) and (A.4), again respectively. We have mentioned that DOP attacks represent a highly sophisticated attacker (see §2.1 and [34]). The principal focus of CFA schemes is on direct control flow attacks, (A.1) and (A.2), rather than data flow attacks, even though the latter can influence the former indirectly. Indeed, some schemes explicitly declare data-flow attacks to be out-of-scope, even though their potential effects on control-flow are recognised [4, 60]. Preventing data-oriented attacks is a burgeoning area of research [95], and it is our belief that designing attestation schemes that detect control- and data-flow attacks is an important direction going forwards.
@@ 9.3 Closed-source Implementations
We also became aware that most CFA proposals are presented without any publicly available artefacts. This is particularly crucial when a scheme demands extensive low-level changes, such as custom CPU extensions as FPGA/HDL modules, binary-level modifications, custom kernel modules, and new assembly routines. A severe cost can be imposed to students and researchers who may wish to rapidly prototype or compare multiple schemes. We also observe that key implementation aspects are sometimes left undefined, such as the method of instrumentation, and some schemes are not fully implemented (6/31, $\approx$20%).
@@ 9.4 On the Need for Common Benchmarks
A wide variety of CFA schemes have been proposed in the scientific literature with a myriad of system architectures and adversarial models. We note that evaluations tend to be conducted in an ad hoc manner, employing different test programs, metrics, and baselines. Consequently, it is difficult to evaluate CFA schemes in order to compare their relative effectiveness. Some proposals have evalauted common programs. In particular, the open syringe pump is the most common, which is an embedded program for remotely controlling liquid injections in medical applications. This has been used by seven CFA proposals [4, 12, 34, 6, 62, 43, 38] that are designed for embedded devices. Using more general benchmarks is another choice in evaluation. SNU Real-Time Benchmark is the most common one evaluated by three proposals [11, 9, 40]. Some proposals [14, 36, 42] published recently opt for Embench-IOT benchmark suite [96]. However, they are not designed for CFA schemes.

### v4

@@ 5.3.2 Trace Modules
ZEKRA [39] implicitly relies on trusted components for measurement and reporting with a certified asymmetric key pair in order to provide path authenticity. The proposal does not specify an implementation per se, but specifies that the module must chronologically and faithfully provide a trace of the executed path when executing a program region to be attested.
@@ 5.3.3 Performance Monitoring Units
LightFAt relies on a CPU’s PMU on $\mathcal{P}$ to monitor the number of instructions per cycles and L1 cache accesses used during the execution of the target program for a given challenge from $\mathcal{V}$. The information about control-flow events is sent to $\mathcal{V}$ who then uses a machine learning model to classify the program execution as being malicious or legitimate. The model is assumed to be trained on the target program pre-deployment with known, labelled malicious/legitimate traces. Results of 95% accuracy and 0.9757 F1-score are presented using a AMD Ryzen 7 2700X 64-bit processor.
5.4 Discrete Modules
@@
Geden & Rasmussen [50] proposed a hardware security module (HSM) that collects control-flow information from the system bus (Figure 14). The HSM continuously takes control-flow events as input from the CPU, including the requested instruction address, opcode, and operand values over the system-on-chip’s control line and address and data buses. The collected events are compared with security policies defined in the HSM’s runtime integrity model for detecting code injection and re-use attacks. The proposal allows $\mathcal{V}$ to acquire a report from the HSM showing whether any security policy violations were detected; it also supports local verification whereby the module outputs a binary result to indicate if a target control-flow hijacking attack is detected. In scenarios where attacks are detected, the module outputs a binary result and provides auxiliary information, such as the last executed instruction.
@@
Tiny-CFA [66] uses a minimalist approach to control-flow attestation, focusing on reducing the trusted computing base and the number of invasive modifications needed to realise a trust anchor. The work employs a simple external hardware module that monitors run-time control-flow states. Tiny-CFA proposes an approach that combines automated code instrumentation with APEX [78], which uses a bespoke hardware module to provide a proof of execution (PoX) from a micro-controller unit core.
@@ 6.1 Compiler-assisted Instrumentation
Where $cp_{A}$ and $cp_{B}$ are checkpoint identifiers representing an arbitrary sequence of BBLs. $H(LoA)$ is the hashed measurement value of a list of actions (LoA), which is the list of CFG edges for traversing $cp_{A}\to cp_{B}$; and $(BBL_{si},BBL_{di})$ is the source and destination BBL at index $i$. The compiler is also used for detecting and instrument control-flow events, which are passed to three kernel-space modules responsible for generating measurements. To support this, the Measurement Generator analyses the LLVM intermediate representation (IR) of the target program’s processed source code. The CRAB framework is used to generate the CFG itself, which is used to resolve indirect forward jumps. We note that mapping LLVM to raw assembly BBLs involves the use of dummy code, which is later removed using a custom binary rewriting tool. The authors add approximately 3.5K lines of code (LOC) to CRAB and LLVM 5.0.
@@
LAPE [52] uses the LLVM-compatible ACES [83] compiler to create and instrument ‘compartments’ on ARMv7-M devices. Programs execute within components where indirect calls and function entry and exit points are recorded and measured using a hash-based approach. We note that, despite LLVM’s flexibility, there are some downsides. LLVM is typically not used unchanged: proposals tend to require extensive modifications or extensions of the magnitude of 1,000s LOC [10, 13, 14]. In practice, this may pose compatibility issues when adapting the schemes to existing toolchains and development environments.
In other work, Hu et al. [11] use a combination of compiler-assisted instrumentation and binary rewriting—discussed in §6.2—to hook control-flow transfer instructions. The GCC compiler with the -finstrument-functions flag is used to instrument function entry and exit points. Afterwards, branches within functions are hooked using binary rewriting using the Capstone disassembly framework [84].
@@ 6.2 Binary Analysis and Rewriting
Binary rewriting involves making direct modifications to the compiled target program, where control-flow transfer instructions are represented and changed in the form of machine code. This approach can be applied to existing programs, making it suitable for incorporating CFA into software where the original source code is not accessible. The approach has particular applications to retrospectively protecting legacy and proprietary software.
@@
Recording every function call can impose significant performance overhead (see p. 3, [62]). To address this, ReCFA [62] derives a CFG from the target program using TypeArmor [85] for performing call-site filtering, whereby CFG nodes are skipped where none of its predecessors has more than one successor. ReCFA uses Dyninst [86] for the static binary instrumentation of user-space programs on $\mathcal{P}$. Dyninst provides a set of libraries and utilities designed for the static (and dynamic) inspection and manipulation of executables. Dyninst enables users to insert, modify, or remove code in a program’s CFG without access to its source code. Intel’s MPK system is used to protect the integrity of instrumented data structures—the loop stack and path stack—used by control-flow events. The authors compare this to SCaRR [13] where control-flow edges are measured in kernel space, which has a “potential enormous context switching cost for the control-flow events on an order of magnitude in gigabytes” (p. 6, [62]). The authors combine kernel mode-based functions to preserve code integrity and MPK-protect data structures in order to uphold the integrity of the measurement procedure.
@@ 7.2 Other Path Log Representations
Another example is RAGE [45], which relies on a machine learning-based approach using unsupervised graph neural networks to avoid computing a complete CFG of the target program. In the training phase, $\mathcal{V}$ uses one legitimate control-flow path represented as a sequence of BBL addresses as a partial CFG (PCFG) with which feature extraction is performed. The features are used to train a Variational Graph Autoencoder (VGAE) model [93] that encodes CFGs into a lower-dimensional space (graph embedding). The legitimate control-flow path is converted to an embedding which serves as a reference. Ten legitimate control-flow paths are used to set an appropriate threshold using the directed Hausdorff distance [94] to quantify the difference between two embeddings. During the attestation process, $\mathcal{P}$ sends a measurement result, a sequence of BBL addresses, executed by the target program to $\mathcal{V}$. $\mathcal{V}$ inputs this data to the VGAE model to produce a new embedding. $\mathcal{V}$ compares the embedding with the referenced embedding; if their distance exceeds a predefined threshold, then it indicates the detection of attacks. The authors tested their proposal using synthetic ROP and DOP attacks, yielding results of 0.9803 (ROP) and 0.9101 (DOP) F1-score while maintaining a low False Positive Rate of 0.0319.
@@ 7.5 Hybrid Measurement
We define hybrid measurement as combining more than one of the measurement methods above. Currently, three existing CFA proposals fall into the category, which combine hash-based measurement and abstract execution. LO-FAT [12] uses abstract execution for loop paths and hash-based measurement for the rest of the control-flow path; avoiding hash-based measurements for loops avoids the path explosion problem. OAT [6] is the first CFA scheme to explicitly address control-flow and data-only attacks on embedded systems. OAT addresses this through a combination of hash-based measurements and compact execution traces in a TEE (ARM TrustZone).
@@ 8 Evaluation
G&R [50] 2019 Hardware Module $\blacktriangle$ $\blacktriangle$ LVRR (C.2)(C.3)(C.4)(C.7) (A.1)(MP)(A.2)(A.3)(A.4)
Liu et al. [7] 2019 TEE (ARM TrustZone) + PUF Whitelist Binary Rewriting Conventional (C.2)(C.5)(C.6)(C.7) (A.1)(A.2)(A.3)
@@
ReCFA [62] 2021 Kernel Whitelist Binary Rewriting Conventional (C.2)(C.4)(C.5)(C.6)(C.7) (A.1)(MP)(A.2)
@@
C-FLAT Linux [42] 2022 Kernel Hash-based Binary Rewriting Conventional (C.1) (A.1)(A.2)(A.3)
@@
GuaranTEE [10] 2022 TEE (Intel SGX) Abstract Execution Compiler (LLVM) LVRR (C.4)(C.5)(C.6)(C.7) (A.1)(A.2)
@@
ARI [46] 2023 TEE* Hybrid Compiler (LLVM) Conventional (C.8) (A.1$\ddagger$(A.2$\ddagger$(A.3$\ddagger$
@@
ZEKRA [39] 2023 Hardware Module* Hash-based $\blacktriangle$ * VC (C.1) (A.1)(A.2)
RAGE [45] 2024 TEE* Path Log Dynamic (DynamoRIO) Conventional (C.1) (A.2)$\pm$(A.4)$\pm$
LightFAt [5] 2024 PMU Path Log Binary Rewriting Conventional (C.1) (A.1)$\pm$±$\ddagger$(A.2)$\pm$±$\ddagger$
@@ 8.1 Comparison Criteria
  • $\mathcal{P}$-$\mathcal{V}$ paradigm: We give the prover-verifier ($\mathcal{P}$-$\mathcal{V}$) paradigm used by the proposal; for example, whether it uses a conventional remote attestation or another approach described in §4.
@@ 8.2 Discussion and Analysis
Many CFA schemes claimed the ability to detect code injection attacks, i.e. (A.1). However, a significant minority of proposals (7/31, $\approx$23%) do not provide this due to features bestowed by the scheme itself. Rather, they make assumptions regarding the availability of mechanisms that enforce some aspect of the target program’s immutability; for example, using executable-space, data execution prevention, and other memory protection mechanisms [12, 50]. LO-FAT, for instance, assumes that “the adversary cannot modify program code at run-time (marked as rx)” (p. 3, [12]) as an implicit requirement for a memory protection mechanism. Geden & Rasmussen [50] requires non-writable code regions and non-executable data regions. In a TEE-based scheme, OAT [6] assumes that the attackers cannot tamper with the instrumented code or the trampoline library in the normal world, citing some memory protection methods for embedded devices as potential solutions [97, 98]. [^8] We note that, in some cases, it is possible to bypass memory protection mechanisms, e.g. DEP, using ROP-style attacks. To this end, we refer the reader to work by Göktas et al. [99].
We observe that ROP attacks are the primary threat in schemes that address code reuse attacks, (A.2). Such proposals add detection mechanisms to control-flow transfers that return from functions. Some work, e.g. C-FLAT [4], address a particular variant in which a function is invoked from multiple calling locations. Suppose there are two legitimate control-flow paths, $A\to B\to C$ and $D\to B\to E$, where $B$ is a vulnerable function. An attacker may perform an attack to alter the control-flow path to $D\to B\to C$, which still visits legitimate CFG nodes and may not be detected as an illicit path. This has been addressed through the use of shadow stacks [50, 62, 7, 13]. When a call-based control-flow transfer is executed, if its target address is valid, then its call-after point will be pushed on this shadow stack. For a return-based transfer, if its target address is valid based on CFG, then $\mathcal{V}$ checks if it is identical to the top element of the shadow stack. If it is not, a control-flow hijacking attack would be detected. If it matches, the top element is popped from the stack. Call-return matching is another solution in which $\mathcal{V}$ marks legitimate pairs of call and return control-flow transfers within the scheme’s offline phase, which are measured at run-time [4].
@@
A final point of note is the relative lack of focus on data-oriented programming (DOP) attacks ((A.4)). Only 5/31 ($\approx$16%) proposals address (A.4). We have mentioned that DOP attacks represent a highly sophisticated attacker (see §2.1 and [38]). The principal focus of CFA schemes is on direct control flow attacks, (A.1) and (A.2), rather than data-oriented attacks, even though the latter can influence the former indirectly. Indeed, some schemes explicitly declare data-flow attacks to be out-of-scope, even though their potential effects on control-flow are recognised [4, 64]. Preventing data-oriented attacks is a burgeoning area of research [100], and it is our belief that designing attestation schemes that detect control- and data-flow attacks is an important direction going forwards.
@@ 9.3 Closed-source Implementations
We also became aware that most CFA proposals are presented without any publicly available artefacts. Only 8/31 ($\approx$26%) of CFA proposals make their implementation publicly available. This is particularly crucial when a scheme demands extensive low-level changes, such as custom CPU extensions as FPGA/HDL modules, binary-level modifications, custom kernel modules, and new assembly routines. This can impose a severe cost to students and researchers who may wish to rapidly prototype or compare multiple schemes. We also observe that key implementation aspects are sometimes left undefined, such as the method of instrumentation, and some schemes are not fully implemented for evaluation (6/31, $\approx$20%).
@@ 9.4 On the Need for Common Benchmarks
A wide variety of CFA schemes have been proposed in the scientific literature with a myriad of system architectures and adversarial models. We note that evaluations tend to be conducted in an ad hoc manner, employing different test programs, metrics, and baselines. Consequently, it is difficult to evaluate CFA schemes in order to compare their relative effectiveness. Some proposals have evaluated common programs. In particular, the open syringe pump is the most common, which is an embedded program for remotely controlling liquid injections in medical applications. This has been used by seven CFA proposals [4, 12, 38, 6, 66, 46, 41] that are designed for embedded devices. Using more general benchmarks is another choice in evaluation. SNU Real-Time Benchmark is the most common one evaluated by three proposals [11, 9, 43]. Some proposals [14, 39, 45] published recently opt for Embench-IOT benchmark suite [101]. However, they are not designed for CFA schemes.